When conducting basic market analysis, we often assume that we have a perfectly competitive market. This setting requires a number of conditions. Which of the following is not a required condition?

  • A. There are many buyers.

  • B. Market participants (buyers and sellers) know what other participants are doing.

  • C. There are many sellers.

  • D. There is no barrier for buyers and sellers to enter or exit the market.

  • E. Some sellers are competitive enough to become dominant players in the market.

Ch2

The following is the labor productivity table for firms A and B, given that they have a fixed number of labor hours. Please complete the opportunity cost table and identify absolute advantages and comparative advantages.

Productivity
Product X Product Y
Firm A 40 units per hour 50 units per hour
Firm B 60 units per hour 100 units per hour

In terms of producing X, which firm has the absolute advantage?   (Enter 1 for firm A or 2 for firm B)

In terms of producing Y, which firm has the absolute advantage?   (Enter 1 for firm A or 2 for firm B)

Opportunity Cost Table
Product X Product Y
Firm A   units of Y per X   units of X per Y
Firm B   units of Y per X   units of X per Y

In terms of producing X, which firm has the comparative advantage?   (Enter 1 for firm A or 2 for firm B)

In terms of producing Y, which firm has the comparative advantage?   (Enter 1 for firm A or 2 for firm B)
Answer #1

From the given information, the statement some sellers are competitive enough to ...players in the market is not the assumpProduct X Product Y 50/40 = 1.25 units of Y per 40/50 = 0.8 units of Xper Y Firm A Firm B 100/60= 1.67 units of Y per 60/100=